OpenStudy (anonymous):

What is the slope of a line with equation y =1/2x+3.

OpenStudy (anonymous):

1/2 the line is in the form y=mx+b, where m is slope and b is the y intercept.

OpenStudy (anonymous):

y=mx +c so the slope is m

OpenStudy (anonymous):

y=mx+c who taught you that?

OpenStudy (anonymous):

it could be b or c just by looking at the form of the equation

OpenStudy (anonymous):

ok just never seen it put that way

OpenStudy (anonymous):

How do i solve this?

Parth (parthkohli):

Do you know what "coefficient" means? @DustinMHayes

OpenStudy (anonymous):

you dont solve anything. its a standard formula. you recognize the formula and plug in what you know

OpenStudy (anonymous):

you just look at the form of the equation your slope is m=1/2 right?

OpenStudy (anonymous):

Y= (1/2)X+3. Is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the only thing that you are looking for is the (1/2)